# FMGS: Foundation Model Embedded 3D Gaussian Splatting for Holistic 3D Scene Understanding

### v1

@@ Abstract
Precisely perceiving the geometric and semantic properties of real-world 3D objects is crucial for the continued evolution of augmented reality and robotic applications. To this end, we present Foundation Model Embedded Gaussian Splatting (FMGS), which incorporates vision-language embeddings of foundation models into 3D Gaussian Splatting (GS). The key contribution of this work is an efficient method to reconstruct and represent 3D vision-language models. This is achieved by distilling feature maps generated from image-based foundation models into those rendered from our 3D model. To ensure high-quality rendering and fast training, we introduce a novel scene representation by integrating strengths from both GS and multi-resolution hash encodings (MHE). Our effective training procedure also introduces a pixel alignment loss that makes the rendered feature distance of same semantic entities close, following the pixel-level semantic boundaries. Our results demonstrate remarkable multi-view semantic consistency, facilitating diverse downstream tasks, beating state-of-the-art methods by $\mathbf{10.2}$ percent on open-vocabulary language-based object detection, despite that we are $\mathbf{851\times}$ faster for inference. This research explores the intersection of vision, language, and 3D scene representation, paving the way for enhanced scene understanding in uncontrolled real-world environments. We plan to release the code upon paper acceptance.
@@ 1 Introduction
3D scene understanding is a critical task in various computer vision and robotics applications. Yet, most existing methods primarily concentrate on either 3D geometry and appearance estimation [41, 35, 24] or 3D object detection and scene segmentation trained on datasets with closed sets of classes [14, 17, 38]. However, for an intelligent agent to interact smoothly with the physical world, merely understanding a subset of the space characterized by pre-identified labels is insufficient. Inspired by the latest advancements in foundation models (FMs) with impressive language and vision semantics [40, 1], this paper aims to develop a more natural 3D scene representation. It integrates both geometric and open-vocabulary semantic information, facilitating easy querying for downstream tasks.
In this paper, we utilize Gaussian Splatting [24] as backbone for reconstructing 3D geometry and appearance, which has demonstrated superior performance in terms of rendering quality for novel-view image synthesis and training efficiency. To assist open-vocabulary 3D scene understanding, we rely on pre-train 2D vision-language CLIP [40] and lift the corresponding information into 3D by a novel multi-view training procedure. We note that, in research communities, the system that is most similar to us is LEFR [25], which integrates implicit NERF [35] based scene representation and CLIP embeddings. Compared to LERF, our system develops a different architecture, provides a variety of technical contributions ranging from high efficiency to 3D consistent query, and obtains significantly better results (approximately $10.2$ percent in representative key metrics).
@@
  • Novel semantic scene representation: We introduce a novel approach combining 3D Gaussians (parameterized by mean, covariance, opacity, and spherical harmonics) for geometry and appearance representation, with MHE for efficient language embedding. This approach addresses memory constraints in room-scale scenes including millions of 3D Gaussians.
  • Multi-view consistent language embeddings: Our training process utilizes Gaussian-splatting based rendering from multiple views, ensuring consistency across time and space. Language embeddings remain invariant to viewpoints, enforcing local proximity consistency within Gaussian volumes.
@@ 2 Related Works
Finally, point-based methods are originally widely used for directly processing data from depth sensors, for performing geometrical and semantic computer vision tasks [39, 22]. Point-NeRF [51] efficiently combines point cloud and NeRF to achieve impressive fast view synthesis results. Recently, 3D Gaussian splatting has been proposed to model points as 3D Gaussians for scene representation [24], and achieved state-of-the-art novel view synthesis rendering quality. However, in [24], the number of Gaussians used for scene representation can easily surpass one million, which introduces strict memory and computational requirements for downstream use cases.
@@
Our work is close and directly FMGS comparable to LERF [25] in terms of assumptions about information available at training phase and query time. For example, it does not assume a priori knowledge of query categories at training time which is assumed 3D-OVS [30].
@@ 3.2 Multi-resolution Hash Encoding
Representing a 3D feature field can have many forms. A naive method is to attach a feature vector (or multiple) to each Gaussian, which can be optimized along with other Gaussian parameters (position, covariance, and so on). However, this is extremely costly in terms of computational cost and memory consumption especially when a large number of Gaussians are generated for scene representation. In fact, adding a $512\times 1$ feature vector per Gaussian will increase the number of optimized parameters to be 9.83×9.83$\times$9.83 × under authentic GS parameterization [24] (10 geometric parameters and 48 spherical harmonic appearance parameters per Gaussian) and 65.0×65.0$\times$65.0 × under simplified GS parameterization [23] (5 geometric parameters and 3 appearance parameters per Gaussian).
@@ 4 Method
Our method, i.e. Foundation Model Embedded Gaussian Splatting (FMGS), leverages strengths of both GS and MHE. We rely on GS for efficient and accurate scene geometry representation and on MHE for representing the scene’s language content in a light-weighted manner. Given a set of input images, we compute their camera poses and 3D sparse visual points using an off-the-shelf structure from motion system, e.g., COLMAP [42]. After that we train GS and acquire 3D Gaussians.
@@ 4.1 Feature Field Architecture
To this end, we parameterize our feature field efficiently using MHE. For a given 3D Gaussian $G(\mathbf{\mathbf{x}})$ with mean position $\mathbf{x}$, we first encode $\mathbf{\mathbf{x}}$ to a feature vector $\mathbf{q}=MHE_{\theta}(\mathbf{\mathbf{x}})$ where $\theta$ is our multi-resolution hash table parameters. We subsequently feed this output into an MLP, which generates our language embedding $\mathbf{f}=MLP^{CLIP}_{\phi}(\mathbf{q})$, with $\mathbf{f}$ belonging to $\mathbb{R}^{D}$. We also normalize $\mathbf{f}$ to make it a unit vector.
@@
Figure 3: FMGS Query pipeline: Top: Given a query view to localize a query, FMGS first renders the dense CLIP feature map. Bottom: given an open-vocabulary query, FMGS generates a relevancy map highlighting the relevant part of the rendered CLIP feature map to the query embedding. The highest relevant is colored as red while the lowest relevant part is colored as blue. Note, for visualization simplicity, we show a single-level MHE in this figure while have used multiple in implementations.
@@ 4.2 Embed the Foundation Models
where $\delta$ is a hyperparameter, which is set to be $1.25$ empirically. Mingyang: please explicitly say whether the above equation uses 2-norm or 1-norm or F-norm As seen in Figure 2 where we use PCA to visualize feature maps like FFD [26], we notice that the target CLIP feature map is not fine-grained enough when embedding similarities of neighboring pixels are considered. This results in poor pixel-alignment gradient signals on Gaussians that are not relevant semantically. On the other hand, DINO [7] features give sharp boundaries between objects [2] in terms of embedding similarity, which can be used for additional regularization.
@@
To transfer the characteristic of DINO features while maintaining the CLIP embedding semantics, we (a) add a DINO feature field loss and (b) define a pixel-alignment loss between the DINO and CLIP feature fields. The DINO feature field shares the same hash grid parameters as CLIP and gives the same encoding $\mathbf{q}$ for a given $\mathbf{x}$. Then the DINO feature field outputs $\mathbf{d}=MLP^{DINO}_{\psi}(\mathbf{q})$ where $\psi$ denotes the parameters of the MLP that are not shared with $MLP^{CLIP}_{\phi}$. This feature field is supervised by passing the $sampled\,image$ once to the pre-trained DINO model without scaling, yielding $\mathbf{D}\in\mathbb{R}^{W\times H\times L}$ where $L$ is the DINO feature dimension. We then render $\hat{\mathbf{D}}$ using the same approach as rendering $\hat{F}$. The DINO regularization loss is as follows:
@@
where $\mathcal{P}$ denotes the set of all the pixels in the image, and $\mathcal{N}(i)$ is the $K\times K$ patch kernel around the rendered feature at pixel $i$. This makes the rendered CLIP feature follow the same similarity pattern as the DINO feature. Note that we stop the gradient back-propagation through DINO features in this training loss, which means $MLP^{DINO}_{\psi}$ would not be affected by this loss.
@@
$\mathcal{L}_{total}=\mathcal{L}_{CLIP}+\lambda\mathcal{L}_{DINO}+\gamma\mathcal{L}_{pixel}$ (9)
We take the mean reduction over all the pixels in the image plane when computing different loss terms. We also empirically find out adding $\mathcal{L}_{pixel}$ in later iterations during training produces the best results. In Figure 4, we provide examples of features extracted from foundation models for training and the rendered features generated by our trained hybrid semantic scene representation. It is evident that the rendered feature maps exhibit higher quality when compared to the features obtained directly from the foundation models, owing to our training process enforces multiple-view consistency.
@@
Figure 5: Effect of dot product similarity (dotpsim) loss. From left to right: RGB image, rendered DINO feature without dotpsim, rendered DINO feature with dotpsim, rendered CLIP without dotpsim, and rendered CLIP feature map with dotpsim. From the CLIP feature maps, we can see that objects can be further distinguished from each other and the background. Differences are highlighted in the red boxes.
@@ 4.3 Relevancy Score
At query time, when provided with a query prompt and a viewing direction, FMGS generates a relevancy map that assigns high scores to semantically relevant locations (see Figure 3). To obtain this relevancy map, we first render the feature map $\mathbf{F}$ using our learned semantic feature field via GS rasterization. Then, we calculate the CLIP embedding $\mathbf{f}_{query}$ corresponding to the query prompt.
To obtain the dense relevancy map, we define a set of canonical phrases with CLIP embeddings $\mathcal{F}_{can}$ following the methodology similar to [25]. Then, we compute pairwise softmax scores based on the cosine similarity between the prompt embedding and $\mathbf{f}_{i,j}$, representing the $\mathbf{F}$ at location $(i,j)$, as well as the canonical embeddings for canonical phrases. We take the minimum value of the softmax over all canonical prompts and deem it the relevancy score $r$:
$r_{i,j}=\min_{n}\frac{\exp(\mathbf{f}_{i,j}^{T}\mathbf{f}_{query})}{\exp(\mathbf{f}_{i,j}^{T}\mathbf{f}_{query})+\exp(\mathbf{f}_{i,j}^{T}\mathbf{f}^{n}_{can})},\mathbf{f}^{n}_{can}\in\mathcal{F}_{can}$
With the above definition, the relevancy score is higher when a query embedding is closer to the rendered feature than the canonical features relatively. We follow [25] and choose the following canonical prompts: “object”, “stuff”, “things”, and “texture”. We also find that these work well for a wide range of queries removing the need for tuning these canonical terms. As depicted in Figure 4, we present representative relevancy maps generated by matching the query embedding with our rendered CLIP feature map and the target CLIP feature map from the foundation model used in our training. It’s evident that the relevancy map derived from our rendered CLIP feature map exhibits finer granularity and higher overall quality.
@@ 4.4 Implementation Details
Our approach employs a hash grid for representing language features, which is notably larger than a typical RGB hash grid. This hash grid comprises 24 layers, spanning resolutions from 16 to 512, and possesses a hash table size of $2^{20}$ with an associated feature dimension of 8. The architecture of the CLIP and DINO MLP models used for $MLP^{CLIP}_{\phi}$ and $MLP^{DINO}_{\psi}$ aligns with that of LERF [24]. Furthermore, we leverage the OpenCLIP [12] ViT-B/16 model, which has undergone training on the LAION-2B dataset. Notably, this model operates with an image pyramid that varies in scale from $0.05$ to $0.5$ of image size, encompassing a total of seven scales for pre-computing CLIP feature pyramid. We take the average pooling of the pre-computed CLIP feature pyramid to get the final hybrid CLIP feature for training our semantic embedding field.
@@
To mitigate GPU memory constraints, we strategically select approximately 10% of the Gaussians based on criteria such as high opacity values and a 2D radius in project Gaussian exceeding 2 pixels in at least one training view. Only these selected Gaussians are involved in the rendering process when we train the semantic embeddings. For optimization, we employ the RAdam optimizer with a weight decay of $10^{-9}$. We incorporate an exponential learning rate scheduler, which spans from an initial value of $5\times 10^{-3}$ and gradually decreases to $4\times 10^{-3}$ over the course of 4.2K training steps (after the initial 30K original GS training steps). In our training regimen, all models initially undergo 2.5K steps without the pixel alignment loss being enabled. These training and testing procedures are executed on an NVIDIA RTX A5000 GPU with 24GB of GPU RAM. The semantic feature field training time with a total of 4.2K steps takes about $1.4$ hours. During training, we use weighting factors to balance the CLIP loss ($\lambda=0.2$) and the pixel-alignment loss ($\gamma$ = $0.01$).

### v2

@@ Abstract
Precisely perceiving the geometric and semantic properties of real-world 3D objects is crucial for the continued evolution of augmented reality and robotic applications. To this end, we present Foundation Model Embedded Gaussian Splatting (FMGS), which incorporates vision-language embeddings of foundation models into 3D Gaussian Splatting (GS). The key contribution of this work is an efficient method to reconstruct and represent 3D vision-language models. This is achieved by distilling feature maps generated from image-based foundation models into those rendered from our 3D model. To ensure high-quality rendering and fast training, we introduce a novel scene representation by integrating strengths from both GS and multi-resolution hash encodings (MHE). Our effective training procedure also introduces a pixel alignment loss that makes the rendered feature distance of same semantic entities close, following the pixel-level semantic boundaries. Our results demonstrate remarkable multi-view semantic consistency, facilitating diverse downstream tasks, beating state-of-the-art methods by $\mathbf{10.2}$ percent on open-vocabulary language-based object detection, despite that we are $\mathbf{851\times}$ faster for inference. This research explores the intersection of vision, language, and 3D scene representation, paving the way for enhanced scene understanding in uncontrolled real-world environments. We plan to release the code on the [project page].
@@ 1 Introduction
3D scene understanding is a critical task in various computer vision and robotics applications. Yet, most existing methods primarily concentrate on either 3D geometry and appearance estimation [45, 38, 26] or 3D object detection and scene segmentation trained on datasets with closed sets of classes [15, 19, 41]. However, for an intelligent agent to interact smoothly with the physical world, merely understanding a subset of the space characterized by pre-identified labels is insufficient. Inspired by the latest advancements in foundation models (FMs) with impressive language and vision semantics [44, 1], this paper endeavors to develop a more natural 3D scene representation supporting open-world visual recognition and understanding. It integrates both geometric and open-vocabulary semantic information, facilitating easy querying for downstream tasks such as object detection and semantic segmentation in open-world scenarios.
In this paper, we utilize Gaussian Splatting [26] as the backbone for reconstructing 3D geometry and appearance, which has demonstrated superior performance in terms of rendering quality for novel-view image synthesis and training efficiency. To assist open-vocabulary 3D scene understanding, we rely on pre-train 2D vision-language CLIP [44] and lift the corresponding information into 3D by a novel multi-view training procedure. We note that, in research communities, the system that is most similar to us is LEFR [27], which integrates implicit NERF [38] based scene representation and CLIP embeddings. Compared to LERF, our system develops a different architecture, provides a variety of technical contributions ranging from high efficiency to 3D consistent query, and obtains significantly better results (approximately $10.2$ percent in representative key metrics).
@@
  • Novel semantic scene representation: We introduce a novel approach combining 3D Gaussians (parameterized by mean, covariance, opacity, and spherical harmonics) for geometry and appearance representation, with MHE for efficient semantic embedding. This approach addresses memory constraints in room-scale scenes including millions of 3D Gaussians.
  • Multi-view consistent language embeddings: Our training process utilizes Gaussian-splatting based rendering from multiple views, ensuring consistency across 3D space in static scenarios. Language embeddings remain invariant to viewpoints, enforcing local proximity consistency within Gaussian volumes.
@@ 2 Related Works
Finally, point-based methods are originally widely used for directly processing data from depth sensors, for performing geometrical and semantic computer vision tasks [42, 24]. Point-NeRF [56] efficiently combines point cloud and NeRF to achieve impressive fast view synthesis results. Recently, 3D Gaussian Splatting (GS) has been proposed to model points as 3D Gaussians for scene representation [26], and achieved state-of-the-art novel view synthesis rendering quality. However, in [26], the number of Gaussians used for scene representation can easily surpass one million, which introduces strict memory and computational requirements for downstream use cases.
@@
Our work is close and directly comparable to LERF [27] in terms of assumptions about information available at training phase and query time. For example, it does not assume a priori knowledge of query categories at training time which is assumed 3D-OVS [33].
Recently, several concurrent works have emerged, investigating the distillation of semantic features into GS scene representations [11, 16]. LEGaussians [49] introduces a method for quantizing high-dimensional concatenated CLIP and DINO features into compact ones to conserve memory, and attach to each individual Gaussian. Langsplat [43] segments images by SAM [28] and then inputs hierarchical semantic segments into CLIP to extract semantic features for the segments. To address memory constraints, Langsplat incorporates a scene-specific language autoencoder to encode CLIP features into lower dimensions, which are also attached to each individual Gaussian. Feature 3DGS [64] distills pixel-aligned features from 2D foundation models, such as SAM [28] and LSeg [30], into GS by associating each Gaussian with a learnable vector. However, this suffers from a loss of semantic understanding capability, particularly for long-tail semantics [27]. Different from all the aforementioned methods, which attach semantic features to each Gaussian, we propose to seamlessly integrate 3D Gaussian scene representation together with MHE for efficient semantic encodings. Notably, our method does not necessitate additional scene-specific quantization or auto-decoder steps, thereby preserving the semantic features’ representation capability from foundation models.
@@ 3.2 Multi-resolution Hash Encoding
Representing a 3D feature field can have many forms. A naive method is to attach a feature vector (or multiple) to each Gaussian, which can be optimized along with other Gaussian parameters (position, covariance, and so on). However, this is extremely costly in terms of computational cost and memory consumption especially when a large number of Gaussians are generated for scene representation. In fact, adding a $512\times 1$ feature vector per Gaussian will increase the number of optimized parameters to be 9.83×9.83$\times$9.83 × under authentic GS parameterization [26] (10 geometric parameters and 48 spherical harmonic appearance parameters per Gaussian) and 65.0×65.0$\times$65.0 × under simplified GS parameterization [25] (4 geometric parameters and 4 appearance parameters per Gaussian).
@@ 4 Method
Our method, i.e. Foundation Model Embedded Gaussian Splatting (FMGS), leverages strengths of both GS and MHE. We rely on GS for efficient and accurate scene geometry representation and on MHE for representing the scene’s language content in a light-weighted manner. Given a set of input images, we compute the corresponding camera poses and 3D sparse visual points using an off-the-shelf structure from motion system, e.g., COLMAP [46]. After that we train GS and acquire 3D Gaussians.
@@ 4.1 Feature Field Architecture
To this end, we parameterize our feature field efficiently using MHE. For a given 3D Gaussian $G(\mathbf{\mathbf{x}})$ with mean position $\mathbf{x}$, we first encode $\mathbf{\mathbf{x}}$ to a feature vector $\mathbf{q}=MHE_{\theta}(\mathbf{\mathbf{x}})$ where $\theta$ is our multi-resolution hash table parameters. We subsequently feed this output into an MLP, which generates our language embedding $\hat{\mathbf{f}}=MLP^{CLIP}_{\phi}(\mathbf{q})$, with $\hat{\mathbf{f}}$ belonging to $\mathbb{R}^{D}$. We also normalize $\hat{\mathbf{f}}$ to make it a unit vector.
@@
Figure 3: FMGS Query pipeline: Top: Given a query view to localize a query, FMGS first renders the dense CLIP feature map. Bottom: given an open-vocabulary query, FMGS generates a relevancy map highlighting the relevant part of the rendered CLIP feature map to the query embedding. The highest relevant is colored as red while the lowest relevant part is colored as blue. Note, for visualization simplicity, we show a single-level MHE in this figure while used multiple in implementations.
@@ 4.2 Embed the Foundation Models
where $\delta$ is a hyperparameter, which is set to be $1.25$ empirically. As seen in Figure 2 where we use PCA to visualize feature maps following FFD [29], we notice that the target CLIP feature map is not fine-grained enough when embedding similarities of neighboring pixels are considered. This results in poor pixel-alignment gradient signals on Gaussians that are not relevant semantically. On the other hand, DINO [7] features give sharp boundaries between objects [2] in terms of embedding similarity, which can be used for additional regularization.
@@
To transfer the characteristic of DINO features while maintaining the CLIP embedding semantics, we (a) add a DINO feature field loss and (b) define a pixel-alignment loss between the DINO and CLIP feature fields. The DINO feature field shares the same hash grid parameters as CLIP and gives the same encoding $\mathbf{q}$ for a given $\mathbf{x}$. Then the DINO feature field outputs $\hat{\mathbf{d}}=MLP^{DINO}_{\psi}(\mathbf{q})$ where $\psi$ denotes the parameters of the MLP that are not shared with $MLP^{CLIP}_{\phi}$. This feature field is supervised by passing the $sampled\,image$ once to the pre-trained DINO model without scaling, yielding $\mathbf{D}\in\mathbb{R}^{W\times H\times L}$ where $L$ is the DINO feature dimension. We then render $\hat{\mathbf{D}}$ using the same approach as rendering $\hat{\mathbf{F}}$. The DINO regularization loss is as follows:
@@
where $\mathcal{P}$ denotes the set of all the pixels in the image, and $\mathcal{N}(i)$ is the $K\times K$ patch kernel around the rendered feature at pixel $i$. This makes the rendered CLIP feature follow the same similarity pattern as the DINO feature. Note that we stop the gradient back-propagation through the rendered DINO features in this training loss, which means $MLP^{DINO}_{\psi}$ would not be affected by this loss. $\mathcal{L}_{pixel}$ is also termed as “dotsim” loss for the rest of the paper, since it is formulated by dot product similarity.
@@
$\mathcal{L}_{total}=\lambda\mathcal{L}_{CLIP}+(1-\lambda)\mathcal{L}_{DINO}+\gamma\mathcal{L}_{pixel}$ (9)
We take the mean reduction over all the pixels in the image plane when computing different loss terms. We also empirically find out adding $\mathcal{L}_{pixel}$ in later iterations during training produces the best results. In Figure 4, we provide examples of features extracted from foundation models for training and the rendered features generated by our trained hybrid semantic scene representation. It is evident that the rendered feature maps exhibit higher quality when compared to the raw feature maps obtained directly from the foundation models, owing to our training process enforces multiple-view consistency.
@@
Figure 5: Effect of dot product similarity (dotpsim) loss. From left to right: RGB image, rendered DINO feature without dotpsim, rendered DINO feature with dotpsim, rendered CLIP without dotpsim, and rendered CLIP feature map with dotpsim. The DINO feature maps do not have significant differences with or without dotpsim. From the CLIP feature maps, we can see that objects can be further distinguished from each other and the background. Differences are highlighted in the red boxes.
@@ 4.3 Relevancy Score
At query time, when provided with a query prompt and a viewing direction, FMGS generates a relevancy map that assigns high scores to semantically relevant locations (see Figure 3). To obtain this relevancy map, we first render the feature map $\hat{\mathbf{F}}$ using our learned semantic feature field via GS rasterization. Then, we calculate the CLIP embedding $\mathbf{f}_{query}$ corresponding to the query prompt.
To obtain the dense relevancy map, we define a set of canonical phrases with CLIP embeddings $\mathcal{F}_{can}$ following the methodology similar to [27]. Then, we compute pairwise softmax scores based on the cosine similarity between the prompt embedding and $\hat{\mathbf{f}}_{i,j}$, representing the $\hat{\mathbf{F}}$ at location $(i,j)$, as well as the canonical embeddings for canonical phrases. We take the minimum value of the softmax over all canonical prompts and deem it the relevancy score $r$:
$r_{i,j}=\min_{n}\frac{\exp(\hat{\mathbf{f}}_{i,j}^{T}\mathbf{f}_{query})}{\exp(\hat{\mathbf{f}}_{i,j}^{T}\mathbf{f}_{query})+\exp(\hat{\mathbf{f}}_{i,j}^{T}\mathbf{f}^{n}_{can})},\mathbf{f}^{n}_{can}\in\mathcal{F}_{can}$
With the above definition, the relevancy score is higher when a query embedding is closer to the rendered feature than the canonical features. We follow [27] and choose the following canonical prompts: “object”, “stuff”, “things”, and “texture”. We also find that these work well for a wide range of queries removing the need for tuning these canonical terms. In Figure 4, we present representative relevancy maps generated by matching the query embedding with our rendered CLIP feature map and the target CLIP feature map from the foundation model used in our training. It’s evident that the relevancy map derived from our rendered CLIP feature map overall exhibits finer granularity and higher quality.
@@ 4.4 Implementation Details
Our approach employs a hash grid for representing language features, which is notably larger than a typical RGB hash grid. This hash grid comprises 24 layers, spanning resolutions from 16 to 512, and possesses a hash table size of $2^{20}$ with an associated feature dimension of 8. The architecture of the CLIP and DINO MLP models used for $MLP^{CLIP}_{\phi}$ and $MLP^{DINO}_{\psi}$ aligns with that of LERF [26]. Furthermore, we leverage the OpenCLIP [13] ViT-B/16 model, which has undergone training on the LAION-2B dataset. Notably, this model operates with an image pyramid that varies in scale from $0.05$ to $0.5$ of image size, encompassing a total of seven scales for pre-computing CLIP feature pyramid. The pre-computed feature pyramid is subsequently processed by average pooling to generate the final hybrid CLIP feature for training our semantics embedded field.
@@
To mitigate GPU memory constraints, we strategically select approximately 40% of the Gaussians based on criteria such as high opacity values and a 2D radius of projected Gaussian exceeding 2 pixels in at least one training view. Only these selected Gaussians are involved in the rendering process when we train the semantic embeddings. For optimization, we employ the RAdam optimizer with a weight decay of $10^{-9}$. We incorporate an exponential learning rate scheduler, which spans from an initial value of $5\times 10^{-3}$ and gradually decreases to $4\times 10^{-3}$ over the course of 4.2K training steps (after the initial 30K original GS training steps). In our training regimen, all models initially undergo 2.5K steps without the pixel alignment loss being enabled. These training and testing procedures are executed on an NVIDIA RTX A5000 GPU with 24GB of GPU RAM. The semantic feature field training time with a total of 4.2K steps takes about $1.4$ hours. During training, we use weighting factors to balance the CLIP loss ($\lambda=0.2$) and the pixel-alignment loss ($\gamma$ = $0.01$).